please help -- thank you
1. Determine, algebraically, whether the function f(x) = 3x + 8 and g(x) = *** are inverse
Answer #1

f\left(x\right)=3x+8,\:\:g\left(x\right)=\frac{x-8}{3}

.

\left(f\:\circ \:g\right)\left(x\right)=f\left(g\left(x\right)\right)

\left(f\:\circ \:g\right)\left(x\right)=f\left(\frac{x-8}{3}\right)

\left(f\:\circ \:g\right)\left(x\right)= 3\left(\frac{x-8}{3}\right)+8

\left(f\:\circ \:g\right)\left(x\right)= x-8+8

{\color{Blue} \left(f\:\circ \:g\right)\left(x\right)= x}

.

.

\left(g\:\circ \:f\right)\left(x\right)= g(f(x))

\left(g\:\circ \:f\right)\left(x\right)= g \left ( 3x+8 \right )

\left(g\:\circ \:f\right)\left(x\right)= \frac{3x+8-8}{3}

\left(g\:\circ \:f\right)\left(x\right)= \frac{3x}{3}

{\color{Blue} \left(g\:\circ \:f\right)\left(x\right)= x}

.

here we have {\color{red} \left(f\:\circ \:g\right)(x)=\left(g\:\circ \:f\right)\left(x\right)= x}

so f(x) and g(x) are inverse of each other

.